# Using physics-inspired Singular Learning Theory to understand grokking & other phase transitions in modern neural networks

### v2

@@ 1 Introduction
Classical statistics are insufficient to understand modern machine-learning. For example, well-studied PAC-bounds from Statistical learning theory are vacuous and fail to explain the remarkable generalization power of neural networks. The reason for this comes from a well-kept secret for the last 15 years: neural networks are singular statistical models [wei2022deep]. Namely, unlike regular models, singular models can implement the same function through multiple distinct parameter values. As an example of how this could occur in a singular model, consider that $\forall\alpha>0,\text{ReLU}(x)=\frac{1}{\alpha}\text{ReLU}(\alpha x)$ [hoogland2023weirdtrick]. Moreover, singular models do not exhibit expected statistical behavior. As an example, the Fisher Information matrix, the basis of asymptotic normality theory for MLEs, is often non-invertible at the true parameters in singular models [watanabe2010equations]. Most pressingly, the majority of AI models today (neural networks, LLMs) are singular. Thus, out of a dire need to study them for AI safety [lehalleur2025you] and interpretability, the extremely mathematically rigorous field of singular learning theory (SLT) developed in 2009 has gained renewed attention.
@@
where $\mathcal{F}_{n}$ is the “free energy” of the model, $w_{\alpha}^{*}$ is the $\mathcal{W}_{\alpha}$-global minima of the population negative-log-likelihood, $L_{n}$ is the empirical negative-log-likelihood, and $\lambda_{\alpha}$ is the local learning coefficient (LLC [^3] Note that higher values of $\lambda_{\alpha}$ indicate greater model complexity.) [lau2023local]. From this equation, we gain a principled understanding of the internal model selection a neural network is performing during training [chen2023dynamical]. Specifically, we see that a neural network during training is consistently trying to minimize the free-energy through a tradeoff of the loss $L_{n}(w^{*}_{\alpha})$ and model complexity $\lambda_{\alpha}$.
@@ 3 Experiments
Question 1: Understanding temporality of grokking and phase transitions
@@
Suppose we are training a model, and let us define $\mathcal{W}_{t}$ to be the weights of the model at iteration $t$. Now suppose at iteration $i$ the model has only memorized (i.e. 100% training accuracy but poor testing accuracy) whereas at iteration $j$ the model has grokked (i.e. 100% testing & training accuracy). Let us define $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ to be the model’s free energy, as per the SLT definition, at each of these respective energies. Borrowing from the Arrhenius Reaction Rate Theory in chemical kinetics [arrhenius1889reaktionsgeschwindigkeit], we present the following hypothesis to explain the time it takes to grok:
@@
For our specific experiments, we take $p=53$ and train all our models using a randomly chosen $40\%$ subset of $\mathbb{Z}_{53}\times\mathbb{Z}_{53}$ (remainder $60\%$ was used for validation.) We trained 500 different models and for each of them stored [^5] Time-based metrics were stored across 100 linearly-spaced checkpoints during training. their training loss & accuracy, validation loss & accuracy, and computed pre-grok and post-grok LLCs if grokking occurred. We stored all of this information through use of a Weights & Biases project; in total this process took over 7 days of continuous CPU compute to run.
For our analyses, which we present momentarily, we only considered the 168 (33.6%) runs during which grokking occurred. We paper a histogram of the distribution of observed LLC jumps $\Delta\lambda$ and log times for grokking $\log(r_{i\to j})$ in Figure 2. We paper a plot of $\log(r_{i\to j})$ versus $\Delta\mathcal{F}_{i\to j}$ across our runs in Figure 3.
@@
This relationship observed is consistent with our hypothesis as it is demonstrates a consistent negative slope, albeit with a low $R^{2}$ score.
@@
This experiment is largely similar to the first, except for a few changes. We trained $60$ models (with papering to Weights & Biases) for $4500$ iterations. For each run, we tracked the LLC across 100 linearly & logarithmically-spaced checkpoints. Moreover, for a given run we detected transitions by (1) smoothing out the training loss curve across every $\sim 10$ steps and (2) finding segments in the training loss curve where the loss decreased by $\geq 10\%$ of the total loss decrease throughout training. We observed our end results to vary greatly based on the method we used to detect these transitions. We then computed $\Delta\mathcal{F}_{i\to j}$to be the change in free-energy across any two consecutive transitions. We provide a plot of $\Delta\mathcal{F}_{i\to j}$ versus $\log r_{i\to j}$ with linear fit in Figure 4.
@@
For our second experiment, we choose a problem that a priori has explicit singularities. Consider a 1-layer neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ given by $f(x)=W_{2}W_{1}x$ where $W_{2}\in\mathbb{R}^{d\times r},W_{1}\in\mathbb{R}^{r\times d}$ for constants $r\leq d$. Then $\text{rank}(W_{2}W_{1})\leq r$ and moreover this model is singularity-full as for any invertible $G\in\mathbb{R}^{r\times r}$, $W_{2}W_{1}=(W_{2}G)(G^{-1}W_{1})$. We fix $d:=100$ and use the same methodology as in the previous experiment, where we are now testing LLC $\lambda_{r}$ across 20 choices of rank $r$ linearly spaced from $1$ to $d$. We paper our results in Figure 6.
@@
We can see an extremely strong quadratic fit, peaking when the rank $r=d$. This fit matches the expected results quite nicely. Specifically, it is a standard fact from algebraic geometry that the set $\mathcal{M}_{r}$ of matrices with rank $r$ is a smooth manifold of r(2dr)r(2d-r) dimension. Hence, on that manifold, we can essentially treat our low-rank neural network $f$ as a regular model with r(2dr)r(2d-r) parameters $\implies\lambda_{r}=\frac{1}{2}r(2d-r)\approx-\frac{1}{2}r^{2}+100r$. Note this quadratic matches cleanly with the $0.429r^{2}+87.342r$ terms in our fit.
@@
For our final experiment, we consider an autoencoder $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$. We again fix $d:=100$ and for a given rank $r\leq d$, generate a sample $x\in\mathbb{R}^{d}$ as $x=Az$ where $z\sim\mathcal{N}(0,I_{r})$ and $A\in\mathbb{R}^{d\times r}$. Thus, increasing $r$ increases the dimension of the subspace that our sample $x$ lives in. We design our autoencoder as a sequential concatenation of a ReLU MLP encoder ($d\to 128\to r$ units) and a ReLU MLP Decoder ($r\to 128\to d$) units. We train our autoencoders with MSE and paper our results in Figure 7:

### v3

@@ 1 Introduction
Classical statistics are insufficient to understand modern machine-learning. For example, well-studied PAC-bounds from statistical learning theory are vacuous and fail to explain the remarkable generalization power of neural networks. The reason for this comes from a well-kept secret for the last 15 years: neural networks are singular statistical models [wei2022deep]. Namely, unlike regular models, singular models can implement the same function through multiple distinct parameter values. As an example of how this could occur in a singular model, consider that $\forall\alpha>0,\text{ReLU}(x)=\frac{1}{\alpha}\text{ReLU}(\alpha x)$ [hoogland2023weirdtrick]. Moreover, singular models do not exhibit expected statistical behavior. As an example, the Fisher Information matrix, the basis of asymptotic normality theory for MLEs, is often non-invertible at the true parameters in singular models [watanabe2010equations]. Most pressingly, the majority of AI models today (neural networks, LLMs) are singular. Thus, out of a dire need to study them for AI safety [lehalleur2025you], the extremely mathematically rigorous field of singular learning theory (SLT) developed in 2009 has gained renewed attention.
@@
where $\mathcal{F}_{n}$ is the “free energy” of the model, $w_{\alpha}^{*}$ is a $\mathcal{W}_{\alpha}$-global minima of the population negative-log-likelihood, $L_{n}$ is the empirical negative-log-likelihood, and $\lambda_{\alpha}$ is the local learning coefficient (LLC [^3] Note that higher values of $\lambda_{\alpha}$ indicate greater model complexity.) [lau2023local]. From this equation, we gain a principled understanding of the internal model selection a neural network is performing during training [chen2023dynamical]. Specifically, we see that a neural network during training is consistently trying to minimize the free-energy through a tradeoff of the loss $L_{n}(w^{*}_{\alpha})$ and model complexity $\lambda_{\alpha}$.
@@ 3 Experiments
Question 1: Understanding the temporality of grokking and phase transitions
@@
Suppose we are training a model, and let us define $\mathcal{W}_{t}$ to be the weights of the model at iteration $t$. Now suppose at iteration $i$ the model has only memorized (i.e. 100% training accuracy but poor testing accuracy) whereas at iteration $j$ the model has grokked (i.e. 100% testing & training accuracy). Let us define $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ to be the model’s free energy, as per the SLT definition, at each of these respective iterations. Borrowing from the Arrhenius Reaction Rate Theory in chemical kinetics [arrhenius1889reaktionsgeschwindigkeit], we present the following hypothesis to explain the time it takes to grok:
@@
For our specific experiments, we take $p=53$ and train all our models using a randomly chosen $40\%$ subset of $\mathbb{Z}_{53}\times\mathbb{Z}_{53}$ (remainder $60\%$ was used for validation.) We trained 500 different models and for each of them stored [^5] Time-based metrics were stored across 100 linearly-spaced checkpoints during training. their training loss & accuracy, validation loss & accuracy, and if grokking occurred – computed pre-grok and post-grok LLCs. We stored all of this information through use of a Weights & Biases project; in total this process took over 7 days of continuous CPU compute to run.
For our analyses, which we present momentarily, we only considered the 168 (33.6%) runs during which grokking occurred. We report a histogram of the distribution of observed LLC jumps $\Delta\lambda$ and log times for grokking $\log(r_{i\to j})$ in Figure 2. We report a plot of $\log(r_{i\to j})$ versus $\Delta\mathcal{F}_{i\to j}$ across our runs in Figure 3.
@@
This observed relationship is consistent with our hypothesis as it is demonstrates a consistent negative slope, albeit with a low $R^{2}$ score.
@@
This experiment is largely similar to the first, except for a few changes. We trained $60$ models (with Weights & Biases again) for $4500$ iterations. For each run, we tracked the LLC across 100 linearly & logarithmically-spaced checkpoints. Moreover, for a given run we detected transitions by (1) smoothing out the training loss curve across every $\sim 10$ steps and (2) finding segments in the training loss curve where the loss decreased by $\geq 10\%$ of the total loss decrease throughout training. We observed our end results to vary greatly based on the method we used to detect these transitions. We then computed $\Delta\mathcal{F}_{i\to j}$to be the change in free-energy across any two consecutive transitions. We provide a plot of $\Delta\mathcal{F}_{i\to j}$ versus $\log r_{i\to j}$ with linear fit in Figure 4.
@@
For our second experiment, we choose a problem that a priori has explicit singularities. Consider a 1-layer neural network $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ given by $f(x)=W_{2}W_{1}x$ where $W_{2}\in\mathbb{R}^{d\times r},W_{1}\in\mathbb{R}^{r\times d}$ for constants $r\leq d$. Then $\text{rank}(W_{2}W_{1})\leq r$ and moreover this model is singularity-full as for any invertible $G\in\mathbb{R}^{r\times r}$, $W_{2}W_{1}=(W_{2}G)(G^{-1}W_{1})$. We fix $d:=100$ and use the same methodology as in the previous experiment, where we are now testing LLC $\lambda_{r}$ across 20 choices of rank $r$ linearly spaced from $1$ to $d$. We report our results in Figure 6.
@@
We can see an extremely strong quadratic fit, peaking when the rank $r=d$. This fit matches the expected results quite nicely. Specifically, it is a standard fact from algebraic geometry that the set $\mathcal{M}_{r}$ of matrices with rank $r$ is a smooth manifold of dimension r(2dr)r(2d-r). Hence, on that manifold, we can essentially treat our low-rank neural network $f$ as a regular model with r(2dr)r(2d-r) parameters $\implies\lambda_{r}=\frac{1}{2}r(2d-r)\approx-\frac{1}{2}r^{2}+100r$. Note this quadratic matches cleanly with the $0.429r^{2}+87.342r$ terms in our fit.
@@
For our final experiment, we consider an autoencoder $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$. We again fix $d:=100$ and for a given rank $r\leq d$, generate a sample $x\in\mathbb{R}^{d}$ as $x=Az$ where $z\sim\mathcal{N}(0,I_{r})$ and $A\in\mathbb{R}^{d\times r}$. Thus, increasing $r$ increases the dimension of the subspace that our sample $x$ lives in. We design our autoencoder as a sequential concatenation of a ReLU MLP encoder ($d\to 128\to r$ units) and a ReLU MLP Decoder ($r\to 128\to d$) units. We train our autoencoders with MSE and present our results in Figure 7: